log3(1/9) = -2
Change each logarithmic equation to an equivalent
equation involving an exponent.

Answer by Edwin McCravy(20086)

log3(1/9) = -2
Change each logarithmic equation to an equivalent 
equation involving an exponent.

Just learn the rule

The logarithm equation

logBA = C 

is equivalent to the exponential equation

BC = A

So, using the above rule,

The logarithm equation

log3(1/9) = -2 

is equivalent to the exponential equation

3-2 = 1/9

Edwin